Virus Stopper Inc., a supplier of computer safeguard systems, uses a cost of capital of 10 percent to evaluate average-risk projects, and it adds or subtracts 2 percentage points to evaluate projects of more or less risk. Currently, two mutually exclusive projects are under consideration. Both have a cost of $ 304 and will last 4 years. Project A, a riskier-than-average project, will produce annual end of year cash flows of $ 90 . Project B, of less than average risk, will produce cash flows of $ 298 at the end of Years 3 and 4 only. To the nearest .01, list the NPV of the higher NPV project. Note, if the NPV is negative, place a - sign in front of your answer. Do not use the $ symbol.
Answer #1

NPV = Present Value of inflows - Present Value of outflows

Discount rate for Project A (Riskier Project) = 10 + 2 = 12%

Discount rate for Project B (Less risky project) = 10 - 2 = 8%

NPV of Project A @ 12% is:

= [$90 x {(1 - 1.12-4) / 0.12] - $304

= [$90 x 3.0373] - $304 = $273.36 - $304 = -$30.64

NPV of Project B @ 8% is:

= [$298/1.083] + [$298/1.084] - $304 = $236.56 + $219.04 - $304 = $151.60

The Project with higher NPV is Project B with a NPV if $151.60